# GraphXForm: Graph transformer for computer-aided molecular design with application to extraction

### v1

@@ 2 $\textsc{BestObj}\leftarrow f(m_{0})$
4       $\textsc{Sampled}\leftarrow\emptyset$
5       while $|\textsc{Sampled}|<q_{\max}$ do
6             $m\leftarrow$ sample from $\mathcal{M}$ with probability $\pi_{\theta}\left(m\ |\ m_{0},m\notin\textsc{Sampled}\right)$
7             $\textsc{Sampled}\leftarrow\textsc{Sampled}\cup\{m\}$
8             if $f(m)>\textsc{BestObj}$ then
9                   $\textsc{BestObj}\leftarrow f(m)$
10                   if $|\textsc{Sampled}|>q_{\min}$ then
11                         break
12
13
@@ 2.3 Policy network architecture
The policy network takes a molecule as input and predicts a probability distribution over next actions. To achieve this, we use a simplified version of the Graphormer (Ying et al., 2021) architecture, which treats the molecule’s atoms as an unordered sequence of nodes. This sequence is passed through a stack of transformer layers, with the attention mechanism augmented by bonding information. The resulting latent representations of the nodes are then used to predict action distributions. Figure 2 provides an illustration, and we elaborate on the architecture below:
Figure 2: Flow of a molecule through the policy network of our method GraphXForm. a. We consider the alphabet $\Sigma=(\mathrm{C},\mathrm{N},\mathrm{O})$. The molecule’s underlying graph is augmented with a virtual node (indexed with 0) and embedded into the latent space $\mathbb{R}^{d}$. b. The latent sequence of atoms is passed through a stack of ReZero transformer layers, omitting positional encoding. In the multi-head attention, individual attention scores between atoms are biased with learnable scalars that depend on their bond order. These bias terms are learnable for each transformer layer and attention head individually. c. The sequence output by the transformer is projected through linear layers to generate logits for the distributions $P^{(0)},P^{(1)}$ and $P^{(2)}$.
@@
Following the Graphormer (Ying et al., 2021) model, we introduce a ‘virtual atom’ $a_{0}=0$ into the molecular graph, which is connected to every atom via a special ‘virtual bond’ (see Figure 2). This virtual bond is represented in the bond matrix by an integer outside the range of standard bond orders. The virtual atom functions similarly to the special [CLS]-token in BERT (Devlin, 2018) and accumulates sequence-level information for downstream tasks. Conceptually, the virtual atom acts as an additional message proxy between nodes in the molecular graph.
The sequence $(a_{0},a_{1},\dots,a_{n})$ is embedded into a corresponding sequence of latent node representations $(\bm{\hat{a}}_{0},\bm{\hat{a}}_{1},\dots,\bm{\hat{a}}_{n})\in\mathbb{R}^{d}$, where each $\bm{\hat{a}}_{i}$ is a learnable vector associated with the atom type.
Additionally, let $\bm{z}_{i}\in\mathbb{R}^{d}$ be a learnable embedding that depends on the number of bonds formed by the $i$-th atom. The augmented sequence $(\bm{\hat{a}}_{0},\bm{\hat{a}}_{1}+\bm{z}_{1},\dots,\bm{\hat{a}}_{n}+\bm{z}_{n})$ is then passed through a stack of transformer layers using ReZero normalization (Bachlechner et al., 2021). Importantly, to preserve the permutation equivariance of the network (i.e., the order of the atoms does not matter), no positional encoding is applied.
@@
The stack of transformer layers produces a transformed sequence of node embeddings $(\bm{e}_{0},\bm{e}_{1},\dots,\bm{e}_{n})$, where each $\bm{e}_{i}\in\mathbb{R}^{d}$ corresponds to the $i$-th atom $a_{i}$ in the original sequence. This sequence is used to predict three distributions, $P^{(0)},P^{(1)},P^{(2)}$, which together factorize the full distribution over possible next actions.
Informally, $P^{(0)}$ represents a distribution over whether to terminate the synthesis, add a new atom, or increase the bond order between two existing atoms. The distribution $P^{(1)}$ specifies which atom to select, either as the atom to bond a newly added atom to, or as the first atom of a pair for increasing bond order. Finally, $P^{(2)}$ is used in the case of bond increase to specify the second atom in the pair. Formally, the unnormalized log-probabilities (logits) for these distributions are obtained as follows:
  • $P^{(0)}$: We compute the logits for $P^{(0)}=(p^{(0)}_{0},p^{(0)}_{1},\dots,p^{(0)}_{k},p^{(0)}_{k+1})$ by projecting the vector $\bm{e}_{0}$ using a linear layer $g_{0}\colon\mathbb{R}^{d}\to\mathbb{R}^{k+2}$. Here, $p^{(0)}_{0}$ is the probability of choosing the action DontChange. For $1\leq j\leq k$, $p_{j}$ represents the probability of adding an atom of type $\Sigma_{k}$ (i.e., $\textsc{AddAtom}(j,\cdot)$ without specifying the bonding atom yet). Lastly, $p^{(0)}_{k+1}$ is the probability of increasing the bond between two atoms ($\textsc{IncBond}(\cdot,\cdot)$) without specifying the atom pair.
  • $P^{(1)}$: To obtain $P^{(1)}=(p^{(1)}_{1},\dots,p^{(1)}_{n})$, we apply a linear layer $g_{1}\colon\mathbb{R}^{d}\to\mathbb{R}$ independently to each $\bm{e}_{1},\dots,\bm{e}_{n}$. Then, $p^{(1)}_{j}$ represents the probability of selecting the $j$-th atom to bond to in $\textsc{AddAtom}(\cdot,\cdot)$, or as the first atom in $\textsc{IncBond}(\cdot,\cdot)$.
  • $P^{(2)}$: Similarly, for $P^{(2)}=(p^{(2)}_{1},\dots,p^{(2)}_{n})$, we apply another linear layer $g_{2}\colon\mathbb{R}^{d}\to\mathbb{R}$ to compute the logits for the probability $p^{(2)}_{l}$, representing the likelihood of selecting the $l$-th atom as the second atom in IncBond.
In practice, during inference, we use $P^{(0)}$ and $P^{(1)}$ directly to choose the first atom for both AddAtom and IncBond, and discard $P^{(2)}$. If IncBond has been selected from $P^{(0)}$, we reprocess the molecule through the network, marking the atom chosen from $P^{(1)}$ by adding a distinct learnable embedding to its corresponding sequence element. In this second pass, we use only $P^{(2)}$ to select the second atom for IncBond.
@@
This multi-step action prediction allows us to easily mask actions (i.e., setting their probability to zero in the policy) that would lead to invalid molecules. While checking for actions that would violate constraints adds a small amount of computational overhead, it has a significant benefit: invalid molecules can be immediately disregarded, preventing the network from wasting resources on infeasible designs. Masking invalid actions not only reduces the search space but also speeds up training by avoiding the need for the model to learn through trial and error how to avoid invalid molecules.
@@ 3 Results and Discussion
First, we outline our design tasks, including the objective functions and property prediction methods. Next, we describe the experimental setup common to all methods and outline the specific setup and hyperparameters for our method, followed by a brief overview of the methods used for comparison. We then present preliminary baseline objectives obtained from screening a list of available molecules and discuss our results in relation to those of the other methods.
3.1 Example solvent design tasks
@@
We aim to design solvents for two-phase aqueous-organic systems used in liquid-liquid extraction. Our focus is on two examples motivated by biotechnology, where products are typically produced in an aqueous solution using microorganisms or enzymes. In such processes, products are to be extracted using the organic solvents we aim to design. We assume a spatially uniform temperature of 298 K in both examples.
The first task focuses on the separation of isobutanol (IBA) from water, a common liquid-liquid extraction process. The chosen solvent should be largely immiscible with water (i.e., low solubility exhibited for both the solvent in water and water in the solvent) and possess high affinity for IBA. As is common practice in chemical engineering, we use the partition coefficient $P_{\text{IBA}}^{\infty}$ at small mole fractions of IBA in both phases $x_{\text{IBA}}$:
@@
Our second task centers on an extraction process presented by Peters et al. (2008), who carried out a solvent screening using COSMO-RS as a property predictor. Here, an enzymatic reaction in aqueous medium converts 3,5-dimethoxy-benzaldehyde (DMBA) molecules to (R)-3,3’,5,5’-tetra-methoxy-benzoin (TMB). The task is to find an organic solvent that forms a two-phase system with water. Similar to the IBA task, an optimal solvent should have a high affinity for the product TMB, enabling it to pull TMB out of the aqueous phase. At the same time, however, the solvent should have a low affinity for the educt DMBA. Designing a suitable solvent for this task is extremely challenging because DMBA and TMB possess similar chemical structures and polarities.
@@
Following Peters et al. (2008), we maximize the ratio $P_{\text{TMB}}^{\infty}/P_{\text{DMBA}}^{\infty}$, while additionally enforcing the miscibility gap constraint from Equation 6 leading to the following scalar objective:
@@
To obtain activity coefficients at infinite dilution, we use a state-of-the-art graph neural network (GNN)  (Gilmer et al., 2017; Reiser et al., 2022; Rittig et al., 2023c). Specifically, we employ the Gibbs-Helmholtz (GH-) GNN that was developed by Medina et al. (2023) for predicting activity coefficients at infinite dilution of binary mixtures at varying temperature. The GH-GNN takes the molecular graphs of the two molecules within a binary mixture and the temperature as inputs. First, structural information from the individual molecular graphs and molecular (self-)interactions based on a mixture graph are encoded into a vector representation, the mixture fingerprint. Based on the mixture fingerprint, a multilayer perceptron (MLP) then predicts the parameters of the Gibbs-Helmholtz relationship so that infinite dilution activity coefficients can be predicted with temperature. The GH-GNN is trained in a structure-to-property manner, i.e., it directly learns activity coefficients at infinite dilution as a function of the molecular graphs. A large data set of experimental activity coefficients at infinite dilution from the DECHEMA Chemistry Data Series  (Gmehling et al., 2008) was used for training. For further details on the GH-GNN architecture, we refer to Medina et al. (2023).
@@
We note that many GNN models and other ML models such as transformers have been developed for activity coefficient prediction  (Rittig et al., 2023a; Winter et al., 2022; Damay et al., 2021), also considering the composition-dependency and thermodynamic consistency  (Rittig et al., 2023b; Winter et al., 2023; Rittig & Mitsos, 2024; Specht et al., 2024). We here chose the GH-GNN as it is specialized for activity coefficients at infinite dilution and was trained on a much larger experimental database than the other models, thus covering a larger chemical space  (Medina et al., 2023), which is desirable for molecular design. This model has shown high prediction accuracy, outperforming well-established methods for predicting activity coefficients at infinite dilution such as UNIFAC  (Fredenslund et al., 1975) or COSMO-RS  (Klamt et al., 2010), cf.  (Medina et al., 2023). Future work could investigate further additional activity coefficient models or directly predicting partition coefficients with ML  (Zamora et al., 2023; Nevolianis et al., 2024).
3.2 General computational setup
To ensure comparability across methods in the design task, we restrict the set of available atoms to carbon, nitrogen, and oxygen (alphabet $\Sigma=(\mathrm{C},\mathrm{N},\mathrm{O})$). In the solvent design tasks we consider, molecules are evaluated using objective functions based on a surrogate model for predicting activity coefficients (see Section 3.1). To avoid potential exploitation of the surrogate model by constructing large molecules that are beyond the size of typically used solvents and thus likely to be different from the molecules used to train the surrogate model, we constrain the design to molecules of a moderate size, with a maximum of 25 atoms. The runtime of a method can vary significantly due to differences in implementation or available hardware. One approach to align computational resources is to limit the number of calls to the objective function (Gao et al., 2022). While this method is useful for comparing sample efficiency, it provides limited insight into the overall efficiency of a method when objective function evaluations are relatively inexpensive. Given the efficiency of the surrogate model, objective function evaluations can be parallelized and run in batches very quickly. Therefore, we take a practical approach to align computational resources: each design run of a method is allocated a time budget of eight wall-clock hours on identical hardware, utilizing a single NVIDIA H100 GPU with 80 GB of memory.
3.3 Our method: GraphXForm
Network hyperparameters
For all our experiments, we set the dimension of the latent space $\mathbb{R}^{d}$ to $d=128$ (see Section 2.3). The network consists of eight transformer layers with ReZero normalization, each with eight heads in the multi-head attention, and a feed-forward dimension of 512.
Pretraining
While it is possible for the agent to learn without any prior knowledge of existing compounds, we pretrain the network on known molecules using SMILES strings from the ChEMBL database (Davies et al., 2015). We filtered the database to ensure the molecules contain only atoms within the alphabet $\Sigma$, yielding approximately 370,000 molecules. Each SMILES string is then converted into an action sequence in our graph formulation. Since this conversion is not unique, we arbitrarily select one possible action sequence. The resulting sequences are used to train the model in a self-supervised way, predicting the next action given the previous actions, as outlined in lines 12-14 in Algorithm 1. Training is conducted with a batch size of 128, over a total of 2.5 million batches.
Fine-tuning
Given an objective function $f\colon\mathcal{M}^{\prime}\to\mathbb{R}\cup\{-\infty\}$, we fine-tune the pretrained policy network using the learning algorithm described in Section 2.2. Unless stated otherwise, we initialize the molecule $m_{0}$ as a single carbon atom. We set $s=500$ as the number of top molecules to retain throughout training. In each epoch, we sample between $q_{\text{min}}=10\cdot 1024$ and maximum of $q_{\text{max}}=50\cdot 1024$ unique action sequences. At the end of each epoch, we train the network on 20 batches of size 64, sampled uniformly from the top 500 molecules. We intentionally keep the number of training batches per epoch relatively low for the network’s size to prevent premature overfitting in all considered cases. However, we observe that in most runs, training on more batches does not adversely affect performance and can actually increase the speed of convergence.
3.4 Benchmark methods
We compare our proposed approach with four diverse molecular design methods: Graph GA (Jensen, 2019) (a genetic algorithm (GA) applied to molecular graphs), REINVENT-Transformer (Xu et al., 2024) (a transformer model that generates SMILES strings), Junction Tree VAE (Jin et al., 2018) combined with a GA for optimization in latent molecular space, and STONED (Nigam et al., 2021) (which performs string mutations on the SELFIES (Krenn et al., 2020) representation). For benchmarking, we use the source code provided by Gao et al. (2022). Although these methods employ different paradigms, we apply the same objective function across all methods to guide the design process.
Graph GA and STONED
Graph GA (Jensen, 2019) uses a GA that directly operates on the molecular graph by mutating atoms and fragments, using crossover rules derived from graph matching. It is a strong non-learning method that has been shown to outperform several SMILES-based learning approaches (Jensen, 2019; Gao et al., 2022). In contrast, STONED (Nigam et al., 2021) employs a GA that operates on the string level and manipulates tokens within the SELFIES molecular representation (Krenn et al., 2020).
REINVENT-Transformer
REINVENT-Transformer (Xu et al., 2024) is a recent state-of-the-art method for designing molecules by synthesizing SMILES strings in an autoregressive way. The approach involves a transformer network that is pretrained in a self-supervised manner on known molecules and then fine-tuned with reinforcement learning using a variant of the REINFORCE algorithm (Williams, 1992). REINVENT-Transformer is similar to GraphXForm in that both rely on pretrained transformers and generate molecules autoregressively. However, REINVENT-Transformer constructs molecules in a language-like manner by building a string token by token from an allowed vocabulary, whereas GraphXForm operates directly on the molecular graph by adding atoms and bonds.
This direct manipulation of the molecular graph allows GraphXForm to guarantee the design of feasible molecules by construction, whereas REINVENT-Transformer must learn the SMILES grammar to avoid invalid structures. Additionally, our approach can start from arbitrary structures, unlike the left-to-right string construction in REINVENT-Transformer, making it straightforward to encode specific constraints to limit the search space. For a fair comparison, we pretrain the policy for REINVENT-Transformer using the same dataset as GraphXForm.
Junction Tree VAE

### v2

@@ 2 $\textsc{BestObj}\leftarrow f(m_{0})$
[…]
@@ 2.3 Policy network architecture
The policy network receives a molecule as input and predicts a probability distribution over possible next actions using a simplified version of the Graphormer (Ying et al., 2021). This model treats the molecule’s atoms as an unordered set of nodes and processes them through a stack of transformer layers with the attention mechanism augmented by bonding information. The resulting latent representations of the nodes are then used to predict action distributions.
Splitting actions
To make the network’s predictions more fine-grained, we decompose each action (except DontChange into three sub-actions, each determined by a separate forward pass through the network:
  • Action level 0: The agent decides whether to end the design process (choosing DontChange) or to modify the molecule. In the latter case, it selects the first atom $j$, which can be an already present one for $\textsc{AddBond}(j,l,o)$ or a new one from the alphabet for $\textsc{AddAtom}(j,l,o)$.
  • Action level 1: Given a modification is intended, the agent selects a second atom $l$ from the current molecule, indicating a new bond between $j$ and $l$.
  • Action level 2: Finally, the agent determines the order $o$ of the bond between $j$ and $l$.
Figure 2 provides an illustration, and we elaborate on the architecture below:
Figure 2: Flow of a molecule through the policy network of our method GraphXForm. a. We consider the alphabet $\Sigma=(\mathrm{C},\mathrm{N},\mathrm{O})$. The molecule’s underlying graph is augmented with a virtual node (indexed with 0) and embedded into the latent space $\mathbb{R}^{d}$. Learnable vectors are added to these embeddings to encode the current action level and decisions on previous levels. b. The latent sequence of atoms is passed through a stack of ReZero transformer layers, omitting positional encoding. In the multi-head attention, individual attention scores between atoms are biased with learnable scalars that depend on their bond order. These bias terms are learnable for each transformer layer and attention head individually. c. The sequence output by the transformer is projected through linear layers to generate logits for the distributions $P^{(0)},P^{(1)}$ and $P^{(2)}$.
@@
Following the Graphormer (Ying et al., 2021) model, we introduce a ‘virtual atom’ $a_{0}=0$ into the molecular graph, which is connected to every atom via a special ‘virtual bond’ (see Figure 2). This virtual bond is represented in the bond matrix by an integer outside the range of standard bond orders. The virtual atom functions similarly to the special [CLS]-token in BERT (Devlin et al., 2019) and accumulates sequence-level information for downstream tasks. Conceptually, the virtual atom acts as an additional message proxy between nodes in the molecular graph.
The atom sequence $(a_{0},a_{1},\dots,a_{n})$ is first embedded latent representations $(\bm{\hat{a}}_{0},\bm{\hat{a}}_{1},\dots,\bm{\hat{a}}_{n})$ in $\mathbb{R}^{d}$, where each $\bm{\hat{a}}_{i}$ is a learnable vector corresponding to the atom type. We then augment these representations with additional learnable embeddings that encode the current action level and other relevant information.
Specifically, we introduce a learnable vector $\bm{r}\in\mathbb{R}^{d}$ to represent the current action level. For each atom $i\in{1,\dots,n}$, we further include the embeddings $\bm{w}_{i}^{(0)}$ and $\bm{w}_{i}^{(1)}$, which indicate whether the $i$-th atom has been selected at action levels 0 and 1, respectively. Moreover, let $\bm{z}_{i}\in\mathbb{R}^{d}$ be an embedding that reflects the total number of bonds (i.e., the degree) formed by the $i$-th atom. The final augmented sequence
$(\bm{\hat{a}}_{0}+\bm{r},\bm{\hat{a}}_{1}+\bm{z}_{1}+\bm{w}_{1}^{(0)}+\bm{w}_{1}^{(1)},\dots,\bm{\hat{a}}_{n}+\bm{z}_{n}+\bm{w}_{n}^{(0)}+\bm{w}_{n}^{(1)})$ (3)
is then passed through a stack of transformer layers using ReZero normalization (Bachlechner et al., 2021). Importantly, to maintain permutation equivariance — ensuring that the order of the atoms does not affect the outcome — we do not apply any positional encodings.
@@
The stack of transformer layers outputs a sequence of node embeddings $(\bm{e}_{0},\bm{e}_{1},\dots,\bm{e}_{n})$, where each $\bm{e}_{i}\in\mathbb{R}^{d}$ corresponds to the original atom $a_{i}$. These embeddings are then used to simultaneously predict three probability distributions, $P^{(0)}$, $P^{(1)}$, and $P^{(2)}$, each corresponding to one of the action levels. The agent selects an action based on the distribution associated with the current action level.
The unnormalized log-probabilities (logits) for these distributions are computed as follows:
  • Action level 0 ($P^{(0)}$): We denote the logits for $P^{(0)}$ as
    $(p^{(0)}_{0},p^{(0)}_{1},\dots,p^{(0)}_{k},q^{(0)}_{1},\dots,q^{(0)}_{n})\in\mathbb{R}^{k+1+n}.$ (5)
    The first $k+1$ logits, $(p^{(0)}_{0},p^{(0)}_{1},\dots,p^{(0)}_{k})$, are obtained by projecting the vector $\bm{e}_{0}$ through a linear layer $g_{0}\colon\mathbb{R}^{d}\to\mathbb{R}^{k+1}$. Here, $p^{(0)}_{0}$ corresponds to the DontChange action, and for $1\leq j\leq k$, the logit $p_{j}^{(0)}$ corresponds to adding a new atom of type $\Sigma_{j}$ (i.e., serving as the first parameter $j$ in $\textsc{AddAtom}(j,\cdot,\cdot)$. The remaining logits, $q^{(0)}_{1},\dots,q^{(0)}_{n}$, are computed by applying a linear layer $h_{0}\colon\mathbb{R}^{d}\to\mathbb{R}$ independently to each of the embeddings $\bm{e}_{1},\dots,\bm{e}_{n}$. For $1\leq j\leq n$, the logit $q^{(0)}_{j}$ corresponds to selecting the first parameter $j$ for the $\textsc{AddBond}(j,\cdot,\cdot)$ action.
  • Action level 1 ($P^{(1)}$): The logits $(q^{(1)}_{1},\dots,q^{(1)}_{n})$ for this level are obtained by applying a linear layer $h_{1}\colon\mathbb{R}^{d}\to\mathbb{R}$ independently to each of $\bm{e}_{1},\dots,\bm{e}_{n}$. Here, $q^{(1)}_{l}$ represents the choice of the second parameter $l$ for either $\textsc{AddAtom}(j,l,\cdot)$ or $\textsc{AddBond}(j,l,\cdot)$.
  • Action level 2 ($P^{(2)}$): The logits $(p^{(2)}_{1},\dots,p^{(2)}_{y})$, which correspond to the bond order $o$ in either $\textsc{AddAtom}(j,l,o)$ or $\textsc{AddBond}(j,l,o)$, are computed by projecting $\bm{e}_{0}$ using a linear layer $g_{2}\colon\mathbb{R}^{d}\to\mathbb{R}^{y}$, where $y$ is a predefined maximum bond order.
After the agent selects an action at the current action level, the chosen information is fed back into the network by updating the learnable vectors $\bm{r},\bm{w}_{i}^{(0)}$ and $\bm{w}_{i}^{(1)}$. This updated state is then used for the subsequent forward pass when predicting the next action level.
@@
The multi-step action prediction allows us to easily mask actions (i.e., setting their probability to zero in the policy) that would lead to invalid molecules. While checking for actions that would violate constraints adds a small amount of computational overhead, it has a significant benefit: invalid molecules can be immediately disregarded, preventing the network from wasting resources on infeasible designs. Masking invalid actions not only reduces the search space but also speeds up training by avoiding the need for the model to learn through trial and error how to avoid invalid molecules.
@@ 3 Results and Discussion
We begin by outlining the experimental setup and hyperparameters for our method. Our first case study tackles the goal-directed design tasks from the well-established GuacaMol benchmark (Brown et al., 2019). In our second case study, we explore two solvent design tasks by detailing the goals, their underlying objective functions, and the property prediction methods used. To contextualize our results, we also present preliminary baselines obtained by screening available molecules and compare our findings with those of other approaches.
3.1 GraphXForm: Pretraining and fine-tuning
Network hyperparameters
For all case studies and experiments, we set the latent space dimension to $d=512$ (see Section 2.3). The network comprises ten transformer layers with ReZero normalization, each featuring eight attention heads and a feed-forward dimension of 2048.
Alphabet and pretraining
We define the atom alphabet $\Sigma$ to include C, N, O, F, P, S, Cl, Br, and I, along with variants for ionization and chirality (see Supplementary Information for the complete alphabet). Although the agent is capable of learning without prior knowledge, we pretrain the network on known molecules using SMILES strings from the ChEMBL database (Davies et al., 2015). We filtered the database to include only molecules containing atoms from $\Sigma$ and split it into a training set of approximately 1.5 million molecules and a validation dataset of around 70,000 molecules. Each SMILES string is converted into an action sequence in our graph formulation; since the conversion is not unique, we select one possible sequence arbitrarily. These sequences are used to train the model in a self-supervised manner – predicting the next action given previous actions – as outlined in lines 6–8 of Algorithm 1. Training is performed with a dropout rate of 0.1, a batch size of 512, and over a total of 1.5 million batches.
Fine-tuning
Given an objective function $f\colon\mathcal{M}^{\prime}\to\mathbb{R}\cup\{-\infty\}$, we fine-tune the pretrained policy network using the learning algorithm described in Section 2.2. We train only the weights of the final linear layers $g_{0},g_{2},h_{0},h_{1}$. Unless stated otherwise, we initialize the molecule $m_{0}$ as a single carbon atom. We set $s=100$ as the number of top molecules to retain throughout training. For the TASAR sampling procedure, we use a beam width of $\beta=512$ and a step size of $\sigma=12$; that is, after every four added atoms or bonds, TASAR seeks better alternative solutions. At the end of each epoch, we train the network on 20 batches of size 64, sampled uniformly from the top 100 molecules. We intentionally keep the number of training batches per epoch relatively low to prevent premature overfitting, though in most runs increasing the number of batches does not harm performance and may even speed up convergence.
Runtime and computational budget
The runtime of a method can vary significantly due to differences in implementation and available hardware. A common approach to align computational resources is to limit the number of calls to the objective function (Gao et al., 2022). While this is useful for comparing sample efficiency, it offers limited insight into overall efficiency when objective evaluations are inexpensive. In our case studies, objective function evaluations are relatively cheap: For the solvent design tasks, for example, these are computed by evaluating the designed molecule with a surrogate neural network that can process batches in parallel. Therefore, to provide a practical comparison between our method and competing approaches, each design run is executed until convergence or until a maximum wall-clock time of eight hours is reached. We run all experiments with a single NVIDIA H100 GPU with 80 GB of memory.
3.2 Case study 1: Drug design
@@
We evaluate GraphXForm for de novo molecular design using the 20 goal-directed tasks from the GuacaMol benchmark (Brown et al., 2019). These tasks span drug rediscovery, similarity-based design, multi-property optimization (MPO), and scaffold hopping. We selected GuacaMol as an established benchmark because our primary focus in this study is to demonstrate the overall optimization capability of our method – i.e., achieving molecules with high scores. We note, however, that current comprehensive benchmarking also considers factors such as sample efficiency and molecular diversity (Gao et al., 2022; Thomas et al., 2022, 2024). Such aspects could be considered in future applications and extensions of GraphXForm, e.g., by further investigating replay buffers (Guo & Schwaller, 2024a, c) and the TASAR parameters.
Benchmark methods
For the drug design tasks, we compare GraphXForm against Graph GA (Jensen, 2019) and REINVENT-Transformer (Xu et al., 2024).
Graph GA (Jensen, 2019) employs a genetic algorithm that directly operates on the molecular graph, mutating atoms and fragments using crossover rules derived from graph matching. This non-learning method is highly effective at making fine-grained local modifications, and it has been shown to outperform several SMILES-based learning approaches (Jensen, 2019; Gao et al., 2022; Brown et al., 2019).
In contrast, REINVENT-Transformer (Xu et al., 2024) designs molecules by synthesizing SMILES strings in an autoregressive manner. This method uses a transformer network that is pretrained in a self-supervised fashion on known molecules and then fine-tuned with reinforcement learning via a variant of the REINFORCE algorithm (Williams, 1992). We include REINVENT-Transformer in our comparisons because, like GraphXForm, it relies on pretrained transformers and generates molecules autoregressively. However, while REINVENT-Transformer constructs molecules token by token from a predefined vocabulary, GraphXForm operates directly on the molecular graph by adding atoms and bonds.
GuacaMol results
Table 1 presents the scores of the best molecules found for four representative (Gao et al., 2024) tasks: Ranolazine MPO, Perindopril MPO, Sitagliptin MPO, and Scaffold Hop. A complete list of results across all 20 tasks is provided in Supplementary Table 1. For Graph GA, we report the results as originally published by Brown et al. (2019). For REINVENT-Transformer, we conducted experiments using the source code and pretrained model provided by Gao et al. (2022).
GraphXForm outperforms both Graph GA and REINVENT-Transformer in different drug design tasks. As shown in Table 1, GraphXForm finds molecules with significantly higher scores for the three MPO cases compared to Graph GA and REINVENT-Transformer. For the scaffold hop task, GraphXForm achieves the best possible score of 1 similar to Graph GA. Considering the perfomance across all 20 tasks of the GuacaMol benchmark (see Supplementary Table 1), GraphXForm attains a total summed score of 18.227, compared to 17.983 for Graph GA. Furthermore, Supplementary Table 1 demonstrates that our method overall outperforms other classic baselines from the original GuacaMol paper, as well as a recent optimization method that utilizes multiple GPT agents for drug design (Hu et al., 2024). These results underscore the robust molecular optimization capabilities of GraphXForm.
Table 1: Performance of different molecular design methods across four tasks from the goal-directed GuacaMol (Brown et al., 2019) benchmark. We report the objective function evaluation of the best molecule found.
Method Ranolazine MPO Perindopril MPO Sitagliptin MPO Scaffold Hop
Graph GA  (Jensen, 2019) 0.920 0.792 0.891 1.000
REINVENT-Transformer  (Xu et al., 2024) 0.934 0.679 0.735 0.582
GraphXForm (ours) 0.944 0.835 0.965 1.000
3.3 Case study 2: Solvent design
Objectives
To further evaluate GraphXForm beyond already-established drug design benchmarks, we propose a solvent design task for two-phase aqueous-organic systems used in liquid-liquid extraction. Our focus is on two examples motivated by biotechnology, where products are typically produced in an aqueous solution using microorganisms or enzymes. In such processes, products are to be extracted using the organic solvents we aim to design. We assume a spatially uniform temperature of 298 K in both examples.
The first solvent design task focuses on the separation of isobutanol (IBA) from water, a common liquid-liquid extraction process. The chosen solvent should be largely immiscible with water (i.e., low solubility exhibited for both the solvent in water and water in the solvent) and possess high affinity for IBA. As is common practice in chemical engineering, we use the partition coefficient $P_{\text{IBA}}^{\infty}$ at small mole fractions of IBA in both phases $x_{\text{IBA}}$:
@@ 3.3 Case study 2: Solvent design
Our second solvent design task centers on an extraction process presented by Peters et al.  (Peters et al., 2008), who carried out a solvent screening using COSMO-RS as a property predictor. Here, an enzymatic reaction in aqueous medium converts 3,5-dimethoxy-benzaldehyde (DMBA) molecules to (R)-3,3’,5,5’-tetra-methoxy-benzoin (TMB). The task is to find an organic solvent that forms a two-phase system with water. Similar to the IBA task, an optimal solvent should have a high affinity for the product TMB, enabling it to pull TMB out of the aqueous phase. At the same time, however, the solvent should have a low affinity for the educt DMBA. Designing a suitable solvent for this task is extremely challenging because DMBA and TMB possess similar chemical structures and polarities.
@@
Following Peters et al.  (Peters et al., 2008), we maximize the ratio $P_{\text{TMB}}^{\infty}/P_{\text{DMBA}}^{\infty}$, while additionally enforcing the miscibility gap constraint from Equation 8 leading to the following scalar objective:
@@
We note that further thermodynamic properties, e.g., boiling and melting points, and sustainability indicators, such as biodegradability and toxicity, are highly relevant for the practical effectiveness of solvents. Such properties could be considered as part of the objective function or as additional constraints in future work.
@@
To obtain activity coefficients at infinite dilution, we use a state-of-the-art graph neural network (GNN)  (Gilmer et al., 2017; Reiser et al., 2022; Rittig et al., 2023c). Specifically, we employ the Gibbs-Helmholtz (GH-) GNN  (Medina et al., 2023) that was developed by Sanchez Medina et al. for predicting activity coefficients at infinite dilution of binary mixtures at varying temperature. The GH-GNN takes the molecular graphs of the two molecules within a binary mixture and the temperature as inputs. First, structural information from the individual molecular graphs and molecular (self-)interactions based on a mixture graph are encoded into a vector representation, the mixture fingerprint. Based on the mixture fingerprint, a multilayer perceptron (MLP) then predicts the parameters of the Gibbs-Helmholtz relationship so that infinite dilution activity coefficients can be predicted with temperature. The GH-GNN is trained in a structure-to-property manner, i.e., it directly learns activity coefficients at infinite dilution as a function of the molecular graphs. A large data set of experimental activity coefficients at infinite dilution from the DECHEMA Chemistry Data Series  (Gmehling et al., 2008) was used for training. For further details on the GH-GNN architecture, we refer to  (Medina et al., 2023).
@@
We note that many GNN models and other ML models such as transformers have been developed for activity coefficient prediction  (Rittig et al., 2023a; Winter et al., 2022; Damay et al., 2021), also considering the composition-dependency and thermodynamic consistency  (Rittig et al., 2023b; Winter et al., 2023; Rittig & Mitsos, 2024; Specht et al., 2024). We here chose the GH-GNN as it is specialized for activity coefficients at infinite dilution and was trained on a much larger experimental database than the other models, thus covering a larger chemical space  (Medina et al., 2023), which is desirable for molecular design. This model has shown high prediction accuracy, outperforming well-established methods for predicting activity coefficients at infinite dilution such as UNIFAC  (Fredenslund et al., 1975) or COSMO-RS  (Klamt et al., 2010), cf. Medina et al. (2023). Future work could investigate further additional activity coefficient models or directly predicting partition coefficients with ML  (Zamora et al., 2023; Nevolianis et al., 2024).
Alphabet and size constraints
We focus our solvent design task on organic H-C-N-O chemistry, restricting the available alphabet to $\Sigma=(\mathrm{C},\mathrm{N},\mathrm{O})$ without ionization and chirality. This choice emphasizes the primary building blocks and reduces the likelihood of designing unstable and unsustainable solvents by omitting elements more likely to contribute to these issues. Furthermore, to prevent potential exploitation of the surrogate by generating excessively large molecules that exceed the typical size of solvents, we constrain the design to molecules with no more than 25 atoms.
Benchmark methods
In addition to Graph GA and REINVENT-Transformer, we benchmark GraphXForm for solvent design against STONED (Nigam et al., 2021) and the Junction Tree Variational Autoencoder (JT-VAE) (Jin et al., 2018). For these methods, we use the source code provided by Gao et al. (2022). STONED is a simple yet efficient algorithm that employs a GA operating at the string level, manipulating tokens within the SELFIES molecular representation (Krenn et al., 2020).